L has y-intercept (0, -3) and is parallel to the line with equation y=2/3x+1.
Answer by Edwin McCravy(20086)

L has y-intercept (0, -3) and is parallel to 
the line with equation y = (2/3)x + 1.

We compare the equation 

 y = (2/3)x + 1

to the slope-y-intercept formula

 y = mx + b

and we can see that the slope is m = 2/3 
and the y-intercept is (0,b) = (0,1)

A line parallel to that one has the same 
slope, m = 2/3 but has a different 
y-intercept, (0,b) = (0,-3), so its
equation must be

 y = mx + b

 y = (2/3)x - 3

(just like the original line's equation
but with a different value for b.

Edwin